OpenStudy (anonymous):

[(2/x)+(5/y)]/[(2/x)-(5/y)]=1 is this correct?

jhonyy9 (jhonyy9):

2y+5x xy 2y+5x ------ * ------- = --------- xy 2y-5x 2y-5x

jhonyy9 (jhonyy9):

your answer is wrong

OpenStudy (anonymous):

why? it's 2/x + 5/y --------- 2/x - 5/y should not i cancel the 5/y because of the different signs?

jhonyy9 (jhonyy9):

yes like aproximatly (2-4)/(2+4)

OpenStudy (anonymous):

how come i was wrong?

jhonyy9 (jhonyy9):

i dont know

jhonyy9 (jhonyy9):

first you need calcule the fraction from numerator with get the denominator common after that multiply the nominators by how many needed ,chck what you will get and make this all with farctions from denominator and the last step make the calcule from terms what remain

OpenStudy (anonymous):

you said my answer is wrong because i answered 1. but if i cancel -4 and +4 in (2-4)/(2+4), the answer is 2/2 which is equal to 1 right?

jhonyy9 (jhonyy9):

not is looke 2-4=-2 and 2+4=6

jhonyy9 (jhonyy9):

ok ???

OpenStudy (anonymous):

ok. thanks again.

jhonyy9 (jhonyy9):

well if i have helped to you please push the good answer button

jhonyy9 (jhonyy9):

thank you